hartnn (hartnn):

Evaluate: (1 + 1/3^5 + 1/5^5 + 1/7^5 + ...)/(1 + 1/2^5 + 1/3^5 + 1/4^5 + ...)

OpenStudy (anonymous):

31/32

OpenStudy (anonymous):

What a strange question. This is beyond me!

OpenStudy (anonymous):

1+1/2^5 + 1/3^5 + 1/4^5 +... (1/2^5 + 1/4^5 + 1/6^5+...) + (1+1/3^5 + 1/5^5+...) 1/2^5( 1+1/2^5 + 1/3^5+...)+(1+1/3^5+1/5^5+...)

OpenStudy (anonymous):

So, 1+1/2^5 + 1/3^5 + 1/4^5 +...=1/2^5( 1+1/2^5 + 1/3^5+...)+(1+1/3^5+1/5^5+...) (1-1/2^5) (1+1/2^5+1/3^5+1/4^5+..) = (1+1/3^5+1/5^5+...)

OpenStudy (anonymous):

Now, (1 + 1/3^5 + 1/5^5 + 1/7^5 + ...)/(1 + 1/2^5 + 1/3^5 + 1/4^5 + ...) {(1-1/2^5) (1+1/2^5+1/3^5+1/4^5+..) }/ (1 + 1/2^5 + 1/3^5 + 1/4^5 + ...) (1-1/2^5)/1 (2^5-1)/2^5 31/32

OpenStudy (anonymous):

Oh I didn't see the division part.. I was trying to compute something completely different.

hartnn (hartnn):

thanks @sauravshakya ! u made it seem not too difficult.... though wolf told me \(\sum \limits_1^\infty (1/n^5)=\zeta(5)\) so i was wondering what's \(\zeta(x)\).....but with our method, it isn't required anymore....

OpenStudy (anonymous):

Those are Reimann zeta functions that are quite peculiar.. one of them relates prime numbers to pi.